# Towards Understanding Multi-Task Learning (Generalization) of LLMs via Detecting and Exploring Task-Specific Neurons

### v1

@@ Abstract
Yongqi Leng and Deyi Xiong thanks:  Corresponding author College of Intelligence and Computing, Tianjin University, Tianjin, China {lengyq,dyxiong}@tju.edu.cn
@@ 1 Introduction
The advent and development of LLMs have marked an significant milestone in natural language processing (Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023). LLMs perform instruction tuning on a wide range of tasks (Wei et al., 2021), exhibiting superior capabilities across multiple tasks and even being able to generalize to unseen tasks (Sanh et al., 2021). Despite their effectiveness, the multi-task learning mechanisms of LLMs remain as an open question.
@@
In order to answer these questions, we perform neuronal analysis for LLMs. First, we identify neurons that are highly correlated with a given task by the gradient attribution method (Simonyan et al., 2013). Subsequently, we conduct fine-tuning and deactivation experiments on these neurons, to analyze their impact on the performance of the given task. Results of extensive experiments show that task-related neurons are indeed present in LLMs and they are highly correlated with specific tasks. We hence term them as task-specific neurons.
@@ 2 Related Work
With the development of LLMs, neuronal interpretability has gained much attention in recent years (Sajjad et al., 2022; Luo and Specia, 2024). Existing researches include knowledge storage (Dai et al., 2021), task solving (Wang et al., 2022), sentiment analysis (Radford et al., 2017), privacy preservation (Chen et al., 2024a; Wu et al., 2023a), and model editing (Gu et al., 2023). In MLLMs, studies find the existence of language-related neurons and utilize neuronal analysis to reveal the multilingual mechanisms of MLLMs (Tang et al., 2024; Chen et al., 2024b; Zhao et al., 2024), which greatly contributes to the understanding of MLLMs. In contrast, limited studies are conducted on the neuronal analysis in multi-task learning in LLMs. We hence extend this line of research from multilingual learning to multi-task learning.
@@
Wei et al. (2021) find that LLMs have excellent zero-shot performance after multi-task fine-tuning, which motivates further investigation into cross-task generalization in depth (Hupkes et al., 2022; Grosse et al., 2023). Existing studies have shown that model size (Wei et al., 2021), number of tasks (Sanh et al., 2021), and data quality (Zhou et al., 2024) all affect the performance of generalization, which illustrates that generalization is affected by a variety of factors. There are also some studies that aim to improving the generalization ability of LLMs, such as step-by-step instruction tuning (Wu et al., 2023b) and hierarchical curriculum learning training strategy (Huang et al., 2024). In addition to this, Yang et al. (2024) conduct an empirical study to investigate generalization between tasks at a fine-grained level. Compared to the above studies, we focus more on the provenance of the generalization phenomenon after instruction tuning, and we analyze task-specific neurons to interpret generalization.
@@
Consistent with previous works (Ke and Liu, 2022; Wang et al., 2024), we categorize continuous learning methods into three classes. (1) Rehearsal-based methods mitigate forgetting by replaying data from previous tasks (Su et al., 2019). (2) Regularization-based methods add explicit regularization terms so that knowledge of previous tasks is retained during continuous training (Aljundi et al., 2018). (3) Parameter isolation-based methods assign task-specific parameters to new tasks, thereby reducing interference between tasks (Razdaibiedina et al., 2023; Wang et al., 2023). Our proposed NCFT method follows the philosophy of parameter isolation continuous learning, but unlike prior works, we do not need to introduce additional parameters and also consider the correlation between tasks.
@@ 3.1 Identifying Task-Specific Neurons in LLMs
First, we need to clarify what we define as a neuron. Currently, the dominant architecture for LLMs is the auto-regressive transformer, in which the basic modules are multi-head self-attention (MHA) and feed-forward network (FFN). Here, we focus only on FFN, which have been demonstrated to store a large amount of parametric knowledge (Dai et al., 2021).
@@ 3.3 Exploring Task-Specific Neurons to Mitigate Catastrophic Forgetting of LLMs
Through the analysis of neurons, we find that while multi-task learning can effectively handle multiple tasks, it does not necessarily achieve optimal performance on a single task (see Section 5.1). This is due to parameter interference between tasks. Similarly, catastrophic forgetting is also caused by parameter interference between tasks. Inspired by this correlation, we propose that isolating task-specific neuron parameters during continuous training might mitigate catastrophic forgetting. In order to substantiate this, we introduce a neuron-level continuous fine-tuning method aimed at mitigating catastrophic forgetting in continuous learning.
@@ 4.1 Experimental Setup
We tested two open-source models that perform well on multi-tasks, including Llama-2-7b (Touvron et al., 2023) and Bloom-7b1 (Le Scao et al., 2023). The dataset and evaluation metrics can be found in Appendix A.3.
@@ 4.2 Results
To bolster the dependability of task-specific neurons, we conducted additional fine-tuning experiments. As shown in Table 2, the fine-tuning approach to task-specific neurons yields remarkable improvements compared to the approach of fine-tuning randomly selected neurons ($29.9$ vs $1.5$ in classification tasks while $17.2$ vs $2.7$ in generation tasks). These improvements remain consistent across both task categories (classificatiton and generation). The only task where the improvement is not significant is AmazonFood, since it has a good enough zero-shot result. Appendix A.4 presents results for Bloom-7b1, which demonstrate the same trend.
@@ 5 Experiments: Analyzing Task-Specific Neurons to Interpret Generalization
We analyzed task-specific neurons to understand the multi-task learning mechanisms of LLMs. Based on the analytical approach of Section 3.2, we conducted two sets of experiments, qualitative and quantitative, on various training-test combinations listed in Appendix A.5.
@@ 5.1 Proportion of Task-Specific Neurons
We controlled the proportion of fine-tuned task-specific neurons to conduct experiments on the various training-test combinations listed in Appendix A.5. Figure 2 shows results for all training-test combinations. In each subfigure, we focus only on the trend of each color line. Comparisons between different color lines are meaningless because they represent different tasks.
@@
As the proportion of trained task-specific neurons increases, we find a continuous increasing trend for the performance of generalization from the trained classification tasks to other classification tasks (orange line in Figure 2 (a)). Similarly, the performance of generalization from the trained generation tasks to other classification tasks (orange line in Figure 2 (b)) and from the trained generation tasks to other generation tasks (green line in Figure 2 (b)) shows the same trend. The overlap rate of task-specific neurons between the training and test tasks can be found in Appendix A.7, where it becomes evident that as the proportion of trained task-specific neurons increases, the overlap rate also experiences a significant surge. Consequently, one plausible explanation is that the overlap of task-specific neurons contributes to transfer learning between tasks, ultimately resulting in consistently higher generalization performance. However, no generalization is produced from the trained classification tasks to other generation tasks (green line in Figure 2 (a)), and the test results are similar to the zero-shot results in Table 2. The reason for no generalization observed from classification to generation might be that classification tasks are usually easier that generation tasks as they only need to predict a single label. In contrast, generation tasks need to generate consecutive texts that satisfy the task requirements, which is relatively harder. This observation is consistent with that observed by Yang et al. (2024).
Trade-offs between Specialization and Generalization.
The red line in Figure 2 indicates the average of the performance of generalization and specialization. We find that when the proportion of fine-tuned neurons is $70\%$, the best trade-off between generalization and specialization can be achieved in both experimental settings.
@@ 5.2 Parameters of Task-Specific Neurons
We evaluated the similarity of specific neuron parameters for the training and test tasks (see Section 3.2 for the way to calculate the similarity) aiming to conduct a qualitative analysis of generalization provenance. We trained a separate model (full-parameter training) for each of the six training tasks in the training-test combination in Appendix A.5, denoted as $M_{1},\cdots,M_{6}$. We then tested these models on the six out-of-domain test tasks listed in that combination, denoted as $T_{1},\cdots,T_{6}$. In a particular layer, for model $M_{i}$ and test task $T_{j}$, $\boldsymbol{P}^{i}_{i}$ and $\boldsymbol{P}^{i}_{j}$ are used to denote the task-specific neuron parameters of training task $i$ and test task $j$ in $M_{i}$, respectively. Then, we calculated the cosine similarity between $\boldsymbol{P}^{i}_{i}$ and $\boldsymbol{P}^{i}_{j}$. For test task $T_{j}$, testing across the six trained models provides six similarity measures. We computed the average of these similarities and then investigated how this average similarity varies across different layers of the model, aiming to show knowledge transfer to the test task $T_{j}$. Figure 3 illustrates the similarity of the different layers for three different settings.
@@
Figure 3 (a) shows how the parameter similarity across three classification test tasks. We find that at the bottom layer, the similarity remains notably low. When reaching a certain layer depth, similarity starts to gradually increase. Finally, the similarity drops again to the value close to that at the bottom layer. This observation holds for all three classification tasks. This illustrates that a model learns the shared knowledge between tasks only after a certain number of layers. In this aspect, knowledge transfer occurs, thus contributing to generalization. Chatterjee et al. (2024) provide similar findings in cross-task in-context learning to ours, which show that information transfer across tasks occurs only after a certain layer depth is reached. Although their findings are based on in-context learning, in-context learning can be understood as a form of implicit training without parameter updates (Akyürek et al., 2022; Von Oswald et al., 2023). We consider these findings resonate with each other.
Method Order-1 Order-2 Order-3 Avg. Order-4 Order-5 Order-6 Avg.
SeqFT 46.4 47.3 47.5 47.1 35.6 34.8 33.5 34.6
SeqLoRA 53.6 54.8 53.1 53.8 47.9 49.5 45.7 47.7
NCFT 71.3 70.9 71.6 71.3 70.5 68.3 71.2 70.0
W-NCFT 73.7 72.3 73.8 73.3 73.4 70.1 72.6 72.0
Per-Task FT 77.2 77.2 77.2 77.2 84.5 84.5 84.5 84.5
@@
We further investigated the relationship between the similarity of task-specific neuron parameters and generalization performance. For each test task, we used six models. We then calculated the similarity in each model between the specific neuron parameters of that test task and the specific neuron parameters of the training task used by that model. Finally, we calculated the correlation coefficients between these parameter similarities and the predictions of the six models. Experiment results are reported in Appendix A.8, from which we find that the similarity is highly correlated with the generalization performance.
@@ 6.1 Experimental Setup
We used Llama-2-7b as the model for experiments. We used two continuous learning benchmarks, Standard CL Benchmark and Large Number of Tasks Benchmark (Razdaibiedina et al., 2023), and tested different task orders. Details on the datasets and task order can be found in Appendix A.9 and A.10.
Metrics and Baselines
We used continuous learning performance and forgetting rate as evaluation metrics. The way for calculating the two metrics can be found in Appendix A.11. The baselines can be found in Appendix A.12.
@@ 6.2 Results and Analysis
As shown in Table 3, our proposed method achieves significant improvements compared to the baselines across both task benchmarks. Such improvements are consistent across various sequences of tasks, illustrating the effectiveness and robustness of our approach. Additionally, we find that W-NCFT outperforms NCFT, suggesting that weighting different task-specific parameters based on their similarity enhances the performance of continuous learning. Figure 4 illustrates the forgetting rate across eight stages on the Large Number of Tasks benchmark, and we can find that both NCFT and W-NCFT methods substantially mitigate catastrophic forgetting.
@@ A.1 Taylor Expansion
We follow Xie et al. (2021) to provide the proof of Equation 3.
@@ A.3 Datasets and Metrics for Identifying Neurons Experiments
    For classification tasks, we chose three tasks. They are sentiment classification, including AmazonFood (Keung et al., 2020), SST-2 (Socher et al., 2013); paraphrase detection, including QQP (Wang et al., 2018), Paws (Zhang et al., 2019); and natural language inference, including MNLI (Williams et al., 2017), GPTNLI [^1] https://huggingface.co/datasets/pietrolesci/gpt3_nli.
@@
    For generation tasks, we chose three tasks. They are summary generation, including CNN/DailyMail (Hermann et al., 2015), Xsum (Narayan et al., 2018); question generation, including Sciqa (Welbl et al., 2017), Tweetqa (Xiong et al., 2019); and data-to-text generation, including E2E (Dušek et al., 2020), CommonGen (Lin et al., 2019).
@@
A.4 Results of Identifying Neurons Experiments in Bloom-7b1

### v2

@@ Abstract
Yongqi Leng and Deyi Xiong thanks:  Corresponding author. College of Intelligence and Computing, Tianjin University, Tianjin, China {lengyq,dyxiong}@tju.edu.cn
@@ 1 Introduction
The advent and development of LLMs have marked a significant milestone in natural language processing (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023). LLMs perform instruction tuning on a wide range of tasks (Wei et al., 2022), exhibiting superior capabilities across multiple tasks and even being able to generalize to unseen tasks (Sanh et al., 2022). Despite their effectiveness, the multi-task learning mechanisms of LLMs remain as an open question.
@@
In order to answer these questions, we perform neuronal analysis for LLMs. First, we identify neurons that are highly correlated with a given task by the gradient attribution method (Simonyan et al., 2014). Subsequently, we conduct fine-tuning and deactivation experiments on these neurons, to analyze their impact on the performance of the given task. Results of extensive experiments show that task-related neurons are indeed present in LLMs and they are highly correlated with specific tasks. We hence term them as task-specific neurons.
@@ 2 Related Work
With the development of LLMs, neuronal interpretability has gained much attention in recent years (Luo and Specia, 2024; Shen et al., 2023). Existing researches include knowledge storage (Dai et al., 2022), knowledge conflicts mitigation (Shi et al., 2024), task solving (Wang et al., 2022), sentiment analysis (Radford et al., 2017), privacy preservation (Chen et al., 2024a; Wu et al., 2023a, 2024), and model editing (Gu et al., 2023). In MLLMs, studies find the existence of language-related neurons and utilize neuronal analysis to reveal the multilingual mechanisms of MLLMs (Tang et al., 2024; Chen et al., 2024b; Zhao et al., 2024), which greatly contributes to the understanding of MLLMs. In contrast, limited studies are conducted on the neuronal analysis in multi-task learning in LLMs. We hence extend this line of research from multilingual learning to multi-task learning.
@@
Wei et al. (2022) find that LLMs have excellent zero-shot performance after multi-task fine-tuning, which motivates further investigation into cross-task generalization in depth (Hupkes et al., 2022; Grosse et al., 2023). Existing studies have shown that model size (Wei et al., 2022), number of tasks (Sanh et al., 2022), and data quality (Zhou et al., 2023) all affect the performance of generalization, which illustrates that generalization is affected by a variety of factors. There are also some studies that aim to improving the generalization ability of LLMs, such as step-by-step instruction tuning (Wu et al., 2023b) and hierarchical curriculum learning training strategy (Huang et al., 2024). In addition to this, Yang et al. (2024) conduct an empirical study to investigate generalization between tasks at a fine-grained level. Compared to the above studies, we focus more on the provenance of the generalization phenomenon after instruction tuning, and we analyze task-specific neurons to interpret generalization.
@@
Consistent with previous works (Ke and Liu, 2022; Wang et al., 2024), we categorize continuous learning methods into three classes. (1) Rehearsal-based methods mitigate forgetting by replaying data from previous tasks (Su et al., 2020). (2) Regularization-based methods add explicit regularization terms so that knowledge of previous tasks is retained during continuous training (Aljundi et al., 2018). (3) Parameter isolation-based methods assign task-specific parameters to new tasks, thereby reducing interference between tasks (Razdaibiedina et al., 2023; Wang et al., 2023b). Our proposed NCFT method follows the philosophy of parameter isolation continuous learning, but unlike prior works, we do not need to introduce additional parameters and also consider the correlation between tasks.
@@ 3.1 Identifying Task-Specific Neurons in LLMs
First, we need to clarify what we define as a neuron. Currently, the dominant architecture for LLMs is the auto-regressive transformer, in which the basic modules are multi-head self-attention (MHA) and feed-forward network (FFN). Here, we focus only on FFN, which have been demonstrated to store a large amount of parametric knowledge (Dai et al., 2022).
@@ 3.3 Exploring Task-Specific Neurons to Mitigate Catastrophic Forgetting of LLMs
Through the analysis of neurons, we find that while multi-task learning can effectively handle multiple tasks, it does not necessarily achieve optimal performance on a single task (see Section 5.1). This is due to parameter interference between tasks. Similarly, catastrophic forgetting is also caused by parameter interference between tasks (Zhu et al., 2024; Wang et al., 2024, 2023b). Inspired by this correlation, we propose that isolating task-specific neuron parameters during continuous training might mitigate catastrophic forgetting. In order to substantiate this, we introduce a neuron-level continuous fine-tuning method aimed at mitigating catastrophic forgetting in continuous learning.
@@ 4.1 Experimental Setup
We tested two open-source models that perform well on multi-tasks, including Llama-2-7b (Touvron et al., 2023) and Bloom-7b1 (Scao et al., 2022). We tested two main types of tasks: classification and generation, details of the dataset and evaluation metrics can be found in Appendix A.3.
@@ 4.2 Results
To bolster the dependability of task-specific neurons, we conducted additional fine-tuning experiments. As shown in Table 2, the fine-tuning approach to task-specific neurons yields remarkable improvements compared to the approach of fine-tuning randomly selected neurons ($29.9$ vs $1.5$ in classification tasks while $17.2$ vs $2.7$ in generation tasks). These improvements remain consistent across both task categories (classificatiton and generation). The only task where the improvement is not significant is AmazonFood, since it has a good enough zero-shot result. Appendix A.4 presents results for Bloom-7b1, which demonstrate the same trend. Additionally, we show the impact of inactivating or fine-tuning a particular class of task-specific neurons on other tasks in Appendix A.4.
@@
Group Training Tasks ID Test Tasks OOD Test Tasks
(a) Amazon, QQP, MNLI Amazon, QQP, MNLI SST-2, Paws, GPTNLI
Tweetqa, CommonGen, Xsum
(b) Sciqa, E2E, CNN Sciqa, E2E, CNN SST-2, Paws, GPTNLI
Tweetqa, CommonGen, Xsum
@@ 5 Experiments: Analyzing Task-Specific Neurons to Interpret Generalization
We analyzed task-specific neurons to understand the multi-task learning mechanisms of LLMs. Based on the analytical approach of Section 3.2, we conducted two sets of experiments, qualitative and quantitative, on various training-test combinations listed in Table 3.
@@ 5.1 Proportion of Task-Specific Neurons
We controlled the proportion of fine-tuned task-specific neurons to conduct experiments on the various training-test combinations. Figure 2 shows results for all training-test combinations. In each subfigure, we focus only on the trend of each color line. Comparisons between different color lines are meaningless because they represent different tasks.
@@
As the proportion of trained task-specific neurons increases, we find a continuous increasing trend for the performance of generalization from the trained classification tasks to other classification tasks (orange line in Figure 2 (a)). Similarly, the performance of generalization from the trained generation tasks to other classification tasks (orange line in Figure 2 (b)) and from the trained generation tasks to other generation tasks (green line in Figure 2 (b)) shows the same trend. The overlap rate of task-specific neurons between the training and test tasks can be found in Appendix A.6, where it becomes evident that as the proportion of trained task-specific neurons increases, the overlap rate also experiences a significant surge. Consequently, one plausible explanation is that the overlap of task-specific neurons contributes to transfer learning between tasks, ultimately resulting in consistently higher generalization performance. To this end, we conducted ablation experiments in Appendix A.7 to exclude the effect of the variable of the number of trained parameters, and the results support this conclusion. However, no generalization is produced from the trained classification tasks to other generation tasks (green line in Figure 2 (a)), and the test results are similar to the zero-shot results in Table 2. The reason for no generalization observed from classification to generation might be that classification tasks are usually easier than generation tasks as they only need to predict a single label. In contrast, generation tasks need to generate consecutive texts that satisfy the task requirements, which is relatively harder. This observation is consistent with that observed by Yang et al. (2024).
@@ 5.2 Parameters of Task-Specific Neurons
We evaluated the similarity of specific neuron parameters for the training and test tasks (see Section 3.2 for the way to calculate the similarity) aiming to conduct a qualitative analysis of generalization provenance. We trained a separate model (full-parameter training) for each of the six training tasks in the training-test combination in Table 3, denoted as $M_{1},\cdots,M_{6}$. We then tested these models on the six out-of-domain test tasks listed in that combination, denoted as $T_{1},\cdots,T_{6}$. In a particular layer, for model $M_{i}$ and test task $T_{j}$, $\boldsymbol{P}^{i}_{i}$ and $\boldsymbol{P}^{i}_{j}$ are used to denote the task-specific neuron parameters of training task $i$ and test task $j$ in $M_{i}$, respectively. Then, we calculated the cosine similarity between $\boldsymbol{P}^{i}_{i}$ and $\boldsymbol{P}^{i}_{j}$. For test task $T_{j}$, testing across the six trained models provides six similarity measures. We computed the average of these similarities and then investigated how this average similarity varies across different layers of the model, aiming to show knowledge transfer to the test task $T_{j}$. Figure 3 illustrates the similarity of the different layers for three different settings.
Testset SST-2 Paws GPTNLI Tweetqa CommonGen Xsum
r p-value r p-value r p-value r p-value r p-value r p-value
PCCs 0.87 0.02 0.92 0.01 0.79 0.05 0.96 0.00 0.96 0.00 0.97 0.00
SROCC 0.81 0.05 0.77 0.07 0.81 0.05 0.77 0.07 0.83 0.04 0.71 0.11
@@
Figure 3 (a) shows how the parameter similarity across three classification test tasks. We find that at the bottom layer, the similarity remains notably low. When reaching a certain layer depth, similarity starts to gradually increase. Finally, the similarity drops again to the value close to that at the bottom layer. This observation holds for all three classification tasks. This illustrates that a model learns the shared knowledge between tasks only after a certain number of layers. In this aspect, knowledge transfer occurs, thus contributing to generalization. Chatterjee et al. (2024) provide similar findings in cross-task in-context learning to ours, which show that information transfer across tasks occurs only after a certain layer depth is reached. Although their findings are based on in-context learning, in-context learning can be understood as a form of implicit training without parameter updates (Akyürek et al., 2023; von Oswald et al., 2023). We consider these findings resonate with each other.
@@
We further investigated the relationship between the similarity of task-specific neuron parameters and generalization performance. For each test task, we used six models. We then calculated the similarity in each model between the specific neuron parameters of that test task and the specific neuron parameters of the training task used by that model. Finally, we calculated the correlation coefficients between these parameter similarities and the predictions of the six models. As shown in Table 4, we find that the similarity is highly correlated with the generalization performance.
@@ 6.1 Experimental Setup
We used Llama-2-7b as the model for experiments. We used two continuous learning benchmarks, Standard CL Benchmark and Large Number of Tasks Benchmark (Razdaibiedina et al., 2023), and tested different task orders. Details on the datasets and task order can be found in Appendix A.8.
Metrics
We used continuous learning performance and forgetting rate as evaluation metrics. Let $a_{i,j}$ be the testing accuracy of the $i$-th task after training on $j$-th task, and $A_{i}$ denote the testing accuracy after training on task $i$ alone. The evaluation metrics are:
Method Order-1 Order-2 Order-3 Avg. Order-4 Order-5 Order-6 Avg.
SeqFT 46.4 47.3 47.5 47.1 35.6 34.8 33.5 34.6
SeqLoRA 53.6 54.8 53.1 53.8 47.9 49.5 45.7 47.7
EPI 48.1 48.0 49.0 48.4 42.3 41.8 43.6 42.6
O-LoRA 76.8 75.7 75.7 76.1 73.7 69.2 72.0 71.6
NCFT (Ours) 71.3 70.9 71.6 71.3 70.5 68.3 71.2 70.0
W-NCFT (Ours) 73.7 72.3 73.8 73.3 73.4 70.1 72.6 72.0
Per-Task FT 77.2 77.2 77.2 77.2 84.5 84.5 84.5 84.5
    $\mathrm{CL}=\frac{1}{N}\sum_{i=1}^{N}a_{i,N}$ (4)
Baselines
We used the following continual learning techniques as baselines:
  • SeqFT: training the entire model parameters on a sequence of tasks.
  • SeqLoRA: training fixed-size LoRA parameters on a sequence of tasks.
  • EPI (Wang et al., 2023b): allocating a small portion of private parameters and learns them with a shared pre-trained model.
  • O-LoRA (Wang et al., 2023a): learning tasks in different (low-rank) vector subspaces that are kept orthogonal to each other in order to minimize interference.
  • Per-Task FT: training a separate model for each task.
@@ 6.2 Results and Analysis
As shown in Table 5, our proposed method achieves significant improvements compared to the first three baselines across both task benchmarks. Even comparing O-LoRA (Wang et al., 2023a) there is a very small difference, on the first benchmark, our method is inferior to O-LoRA, but we outperform it on the second benchmark. Such improvements are consistent across various sequences of tasks, illustrating the effectiveness and robustness of our approach. Additionally, we find that W-NCFT outperforms NCFT, suggesting that weighting different task-specific parameters based on their similarity enhances the performance of continuous learning. Figure 4 illustrates the forgetting rate across eight stages on the Large Number of Tasks benchmark, and we can find that both NCFT and W-NCFT methods substantially mitigate catastrophic forgetting.
@@ 7 Conclusion
Acknowledgments
The present research was supported by the National Key Research and Development Program of China (Grant No. 2023YFE0116400). We would like to thank the anonymous reviewers for their insightful comments.
@@ A.1 Taylor Expansion
We follow Xie et al. (2021) and Zhu et al. (2024) to provide the proof of Equation 3.
@@ A.3 Datasets and Metrics for Identifying Neurons Experiments
    For classification tasks, we chose three tasks. They are sentiment classification, including AmazonFood (Keung et al., 2020), SST-2 (Socher et al., 2013); paraphrase detection, including QQP (Wang et al., 2019), Paws (Zhang et al., 2019); and natural language inference, including MNLI (Williams et al., 2018), GPTNLI [^1] https://huggingface.co/datasets/pietrolesci/gpt3_nli.
@@
    For generation tasks, we chose three tasks. They are summary generation, including CNN/DailyMail (Hermann et al., 2015), Xsum (Narayan et al., 2018); question generation, including Sciqa (Welbl et al., 2017), Tweetqa (Xiong et al., 2019); and data-to-text generation, including E2E (Dusek et al., 2020), CommonGen (Lin et al., 2020).
@@
A.4 Additional Experiments for Identifying Neurons